OpenStudy (harrisonshark):

What is the equation of a circle whose center is at the origin and whose radius is 9?

OpenStudy (raden):

x^2 + y^2 = r^ just plug in r = 9

OpenStudy (harrisonshark):

thank you

OpenStudy (raden):

opss... i meant x^2 + y^2 = r^2 you're welcome :)